Define an xy coordinate system on the face of an analog clock with the origin at the center. What is the horizontal component of a 3 cm long hour hand at 10:00 am?
1
Expert's answer
2019-08-30T09:25:35-0400


Calculation of the angle between each hour.


"\theta_{div}=\frac{360^{0}}{12div}=30^{0}/div"


From the direction of the positive x-axis until 10am there are 5 divisions, therefore the degrees are


"\theta=5div*30^{0}/div=150^{0}"


Now, the horizontal component is given by


"cos\theta=\frac{r_{x}}{r}\\
r_{x}=r*cos\theta\\"


Numerically evaluating


"r_{x}=3.00cm*cos(150^{0})\\
r_{x}=-2.60cm"


The horizontal component of the long hand at 10:00am is

"\boxed{r_{x}=-2.60cm}"